OpenStudy (anonymous):

Can anyone help please?! If x-1/x = 3+2 under-root 2, find the value of 1/4 (x to the power 3 - 1/x to the power 3) Pls help!

OpenStudy (haseeb96):

@Agrima_Sharma can u draw the equation.

OpenStudy (haseeb96):

just click on draw button then write it.

OpenStudy (anonymous):

\[If x - \frac{ 1 }{ x } = 3 + 2\sqrt{2} , then find the value of: \frac{ 1 }{ 4 } \left( x ^{3} - \frac{ 1 }{ x ^{3} } \right) \]

OpenStudy (anonymous):